2. -8x-3y=69
8x+7y=-65
how do you solve step by step to find the two points

Found 2 solutions by Alan3354, htmentor:
Answer by Alan3354(69443)

-8x-3y=69
8x+7y=-65
---------- Add
4y = -134
y = -67/2
----------
8x + 7(-67/2) = -65
x = 335/2
-----------
What 2 points?

Answer by htmentor(1343)

The easiest way to solve this system is to add the two equations to eliminate x:
-8x + 8x -3y + 7y = 4
4y = 4
y = 1
Now substitute this value for y into either of the two equations. Using equation 1, we have:
-8x - 3(1) = 69
-8x = 72
x = -9
Ans: x = -9, y = 1
Check:
8(-9) + 7(1) = -65